OpenStudy (anonymous):

Pablo bought a bicycle for $368.90. The price Pablo paid included sales tax of 8.5%. Which equation can be used to find the price, p, of the bicycle before tax? A. 0.085p = 368.90 B. 8.5p = 368.90 C. p + 0.085p = 368.90 D. p + 8.5p = 368.90

OpenStudy (anonymous):

I think it maybe c

OpenStudy (is3535):

(xy=x times y just fyi) bought bike for 368.9 this was including sales tax of 8.5% (percent means part out of 100 so 8.5%=8.5/100=0.85/10= 0.085/1 1=0.085) so paid=price+tax paid=368.9 price=p tax=p times 8.5%=0.085 368.9=p+p times 0.085 368.9=p+0.085p or p+0.085p=368.9

OpenStudy (is3535):

c is your answer

OpenStudy (is3535):

medal

OpenStudy (anonymous):

thx u